OpenStudy (anonymous):

a and b are the length of the legs of a right triangle, and c is the hypotenuse. Find the missing length if a=15 and b=39

OpenStudy (anonymous):

a squared + b squared = c2 so 15 X 15 = 225 + 39 X 39= 1521 225 + 1521 = c squared 1746= c squared square root of 1746 = 41.7852 thats your answer.

OpenStudy (anonymous):

\[It's C is: 3\sqrt{194}\] OR 41.79